The Lakeside Coffee Shop morning customer load follows a normal distribution. The mean number of customers is 60 and the standard deviation is 15. If the store gets more than 70 customers, they need to have a fourth staff member working. What is the probability that they will need a fourth person tomorrow? Include evidence of your work by typing out your solution fully.



1
Expert's answer
2021-08-04T14:40:15-0400

Let "X=" the number of customers: "X\sim N(\mu, \sigma^2)."

Given "\mu=60, \sigma=15."


"P(X>70)=1-P(X\leq 70)"

"=1-P(Z\leq \dfrac{70-\mu}{\sigma})=1-P(Z\leq \dfrac{70-60}{15})"

"=1-P(Z\leq 0.6667)\approx0.2525"

The probability that they will need a fourth person tomorrow is 0.2525.